F(X)= X2 - X
G(X) = X-1
WHAT IS F [G(X)]

Answer by Nate(3500)

f[g(x)] = (x - 1)^2 - (x - 1)
f[g(x)] = x^2 - 2x + 1 - x + 1
f[g(x)] = x^2 - 3x + 2